OpenStudy (anonymous):

Solve the system of equations: (4/x)-(3/y)=0 (6/x)+(3/2y)=2

OpenStudy (luigi0210):

Start by eliminating the fractions.

OpenStudy (dumbcow):

i think for this one i would just do a straight substitution solve first equation for x 4/x = 3/y ---> x/4 = y/3 ---> x = 4y/3 plug into 2nd equation 6/(4y/3) + 3/2y = 2 18/4y + 3/2y = 2 multiply everything by "y" 18/4 + 3/2 = 2y y = 3